you have a piece of wood that is 84 inches long. you cut the wood into 3 pieces. the 2nd piece is 8 inches longer than the first piece. the 3rd piece is 16 inches longer than the second piece. write an equation and then solve it to determine the lengths of the 3 pieces.
Answer by lwsshak3(11628)

you have a piece of wood that is 84 inches long. you cut the wood into 3 pieces. the 2nd piece is 8 inches longer than the first piece. the 3rd piece is 16 inches longer than the second piece. write an equation and then solve it to determine the lengths of the 3 pieces.
**
le x=length of 1st piece
x+8=length of 2nd piece
x+8+16=x+24=length of 3rd piece
..
x+(x+8)+(x+24)=84
3x+32=84
3x=84-32=52
x=52/3=17+1/3
x+8=25+1/3
x+24=41+1/3
..
length of 1st piece=17+1/3 in
length of 2nd piece=25+1/3 in
length of 3rd piece=41+1/3 in